OpenStudy (anonymous):

Miranda bought some apples and oranges at the store for fruit salad. Owen bought three times as many apples and twice as many oranges. Miranda has 5 fruits in her bag, but Owen has 12 fruits in his bag. How many apples did Miranda purchase?

OpenStudy (anonymous):

is owens equation 3a + 2o=12?

OpenStudy (inowalst):

Lets use a for apples and o for oranges. 12=3a+2o M=5 O=12 So yes, you're right.

OpenStudy (anonymous):

and then marandas equation would be a+0=5

OpenStudy (inowalst):

Yes.

OpenStudy (anonymous):

but how do i solve it then

OpenStudy (inowalst):

"Miranda bought some apples and oranges at the store for fruit salad. Owen bought three times as many apples and twice as many oranges."

OpenStudy (anonymous):

well its a multiple choice question its ether 1 2 3 5

OpenStudy (inowalst):

@TheSmartOne I'm kinda stuck on this.

OpenStudy (anonymous):

thanks for trying though

OpenStudy (inowalst):

No problem.

TheSmartOne (thesmartone):

Hmmm.

TheSmartOne (thesmartone):

Well first we would need to label what Miranda's fruits are.

OpenStudy (anonymous):

i think i got it but tell me if i wrong or not

OpenStudy (anonymous):

i think the 'a' term is 17 because i have to single out one of the terms

TheSmartOne (thesmartone):

So if Miranda has oranges and apples. And o + a = 5 because she has 5 fruits so Miranda's equation would be o + a = 5

OpenStudy (anonymous):

and 'o' equals 10

OpenStudy (anonymous):

yes

TheSmartOne (thesmartone):

And Owen has 3 times as many oranges and 2 times as many apples And for Owen, the equation would be 3a + 2o = 12

OpenStudy (anonymous):

yea

TheSmartOne (thesmartone):

So your equation was correct :) And now we use those 2 equations and solve for a :)

OpenStudy (anonymous):

but im trying that and im not getting the right answer

TheSmartOne (thesmartone):

3a + 2o = 12 a + o = 5 So those are our two equations. We can solve them in many ways. There are actually 3 ways to solve them-- Elimination, Substitution, and graphing Which method do you know or do you find the esaist? :)

OpenStudy (anonymous):

substitition

TheSmartOne (thesmartone):

So since we are solving for a, lets make this equation a + o = 5 If we move around terms in that so that we get what o = ? we can substitute that back in and solve for a :)

TheSmartOne (thesmartone):

So a + o =5 o = ?

TheSmartOne (thesmartone):

@inowalst You definitely tagged the correct person ;)

OpenStudy (inowalst):

xD Haha, of course I did!

OpenStudy (anonymous):

i am very confused

TheSmartOne (thesmartone):

So \(\sf a + o =5\) Lets subtract a from both sides :) \(\sf a + o - a = 5- a\) What do we get if we simplify that?

OpenStudy (anonymous):

o = -5a?

TheSmartOne (thesmartone):

Actually, no. \(\sf a - a + o = 5 - a\) Which can be simplified to \(\sf o = 5 - a\)

OpenStudy (anonymous):

oh ok

TheSmartOne (thesmartone):

And now we plug that in to our first equation we had which is \(\sf 3a + 2o = 12\)

TheSmartOne (thesmartone):

So since \(\sf o = 5- a\) we plug that into \(\sf 3a +2o =12\) And now just solve for a :) \(\sf 3a +2o =12\) \(\sf 3a +2\color{red}{(5-a)} =12\)

OpenStudy (anonymous):

ohh ok thank you so much! im not completely sure how to give you a medal but yes

TheSmartOne (thesmartone):

\(\bf\huge~~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

OpenStudy (anonymous):

ok cool thank you

TheSmartOne (thesmartone):

And here is how to give a medal :) http://goo.gl/kTF8qu